A coffee merchant has coffee beens that sell for $9 per pound and $12 per pound. The two types are to be mixed to create 100 lb. of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
Answer by ankor@dixie-net.com(22740)

A coffee merchant has coffee beens that sell for $9 per pound and $12 per pound. The two types are to be mixed to create 100 lb. of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
:
Let x = amt of $12 coffee
:
We know the total will be 100 pounds, therefore
(100 - x) = amt of $9 coffee
:
12x + 9(100-x) = 11.25(100)
:
12x + 900 - 9x = 1125
:
12x - 9x = 1125 - 900
:
3x = 225
:
x = 225/3
:
x = 75 lb of $12 coffee
:
100 - 75 = 25 lb of $9 coffee
:
:
Check: 12(75) + 9(25) = 11.25(100)